# Benchmarking Vision Foundation Models for Input Monitoring in Autonomous Driving

### v2

@@ Results
APS OC-SVM GMM NF
AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95
Trained on ImageNet RN50 0.73 0.69 0.70 0.76 0.71 0.62 0.70 0.67 0.78 0.75 0.71 0.63
RN101 0.70 0.64 0.68 0.72 0.65 0.60 0.69 0.65 0.81 0.75 0.70 0.63
RN152 0.68 0.63 0.74 0.72 0.68 0.77 0.67 0.62 0.77 0.70 0.65 0.73
ViT-B/16 0.79 0.75 0.53 0.83 0.81 0.53 0.81 0.78 0.53 0.82 0.79 0.56
ViT-L/16 0.88 0.85 0.39 0.86 0.83 0.47 0.88 0.85 0.40 0.88 0.85 0.42
SwinB 0.84 0.82 0.53 0.83 0.84 0.54 0.85 0.84 0.53 0.86 0.85 0.44
SwinT 0.72 0.64 0.62 0.74 0.70 0.75 0.73 0.66 0.65 0.72 0.67 0.77
Autoencoder 128 0.66 0.61 0.71 0.62 0.56 0.74 0.62 0.55 0.79 0.42 0.43 0.94
256 0.68 0.62 0.69 0.68 0.64 0.77 0.52 0.48 0.87 0.60 0.56 0.82
512 0.65 0.61 0.73 0.62 0.58 0.74 0.61 0.58 0.88 0.55 0.53 0.88
1024 0.70 0.67 0.74 0.62 0.57 0.84 0.65 0.60 0.83 0.68 0.63 0.65
Ours Clip-RN50 0.85 0.82 0.49 0.90 0.88 0.36 0.85 0.83 0.51 0.82 0.78 0.52
Clip-RN50x64 0.94 0.93 0.24 0.94 0.93 0.23 0.94 0.93 0.24 0.88 0.84 0.37
SwinB 0.88 0.84 0.47 0.91 0.89 0.32 0.84 0.80 0.54 0.95 0.94 0.16
@@
Trained on ImageNet RN50 0.90 0.89 0.33 0.94 0.93 0.23 0.92 0.91 0.27 0.97 0.97 0.12
RN101 0.89 0.87 0.36 0.93 0.92 0.27 0.91 0.90 0.30 0.97 0.96 0.12
RN152 0.90 0.89 0.35 0.95 0.94 0.18 0.93 0.91 0.29 0.97 0.96 0.10
ViT-B/16 0.86 0.82 0.38 0.93 0.92 0.24 0.91 0.88 0.26 0.97 0.96 0.12
ViT-L/16 0.88 0.84 0.31 0.92 0.91 0.24 0.91 0.88 0.23 0.98 0.97 0.07
SwinB 0.90 0.88 0.31 0.96 0.96 0.14 0.95 0.93 0.19 0.99 0.99 0.05
SwinT 0.88 0.84 0.31 0.95 0.93 0.16 0.93 0.90 0.20 0.99 0.98 0.04
Autoencoder 128 0.60 0.59 0.87 0.59 0.56 0.86 0.60 0.58 0.88 0.55 0.52 0.87
256 0.60 0.59 0.85 0.58 0.55 0.87 0.54 0.53 0.90 0.59 0.55 0.85
512 0.60 0.60 0.84 0.56 0.55 0.87 0.60 0.59 0.88 0.57 0.54 0.85
1024 0.57 0.55 0.87 0.60 0.57 0.86 0.52 0.51 0.86 0.63 0.59 0.81
Ours Clip-RN50 0.76 0.72 0.59 0.85 0.82 0.43 0.75 0.71 0.64 0.95 0.92 0.16
Clip-RN50x64 0.99 0.99 0.03 1.00 1.00 0.01 1.00 1.00 0.01 0.99 0.99 0.04
SwinB 1.00 1.00 0.02 1.00 1.00 0.00 1.00 0.99 0.02 1.00 1.00 0.00
@@
Table 6: Comparative evaluation of out-of-distribution detection performance using ImageNet-pretrained models, autoencoders, and VFMs on Lost and Found [57] and Bravo-Synrain [49] benchmarks. We evaluate using AUROC, AUPR, and FPR95 metrics across four methods: APS, OC-SVM, GMM, and NF. For VFMs, we utilize image encoders from CLIP [59] with ResNet-50 and ResNet-50x64 backbones, and from Grounding-DINO [44] with SwinB backbone. Best performances for each metric and dataset are highlighted in bold.
A.3 Real-time capability: Computational Complexity of Chosen Methods
@@
A primary consideration for our approach is its operational feasibility in real-time AD systems. Our method leverages only the image encoder component of Vision Foundation Models (VFMs) for feature extraction, specifically employing Swin Transformer backbones (i.e. SwinB and SwinT) integrated within the DETR framework [44], and ResNet variants (e.g. RN50, RN50x64) from CLIP [59]. These are shown to be real-time capable:
@@
  • Recent work by Zhao et al. [89] demonstrates that DETR-based architectures can achieve real-time performance while surpassing YOLO-based models in object detection.
@@
  • Furthermore, Kassab et al. [36] successfully deployed CLIP’s ResNet backbones for real-time visual scene understanding in SLAM applications, validating the computational feasibility of these architectures.
These findings collectively support the practical viability of our approaches in AD deployments. Additionally, our framework offers flexibility in deployment scenarios: the system can be configured to
  • process every $n$th frame or
  • operate on downsampled inputs,
enabling practical trade-offs between computational efficiency and monitoring frequency.
There are differences in the real-time capability of the distribution modeling approaches: In our experiments, we evaluate one similarity-based approach (Average Pairwise Similarity (APS)) and three distribution modeling methods (One-Class Support Vector Machine (OC-SVM), Gaussian Mixture Model (GMM), and Normalizing Flows (NF)). While APS directly computes distances between $\mathcal{X}_{\mathrm{monitor}}$ and $\mathcal{X}_{\mathrm{test}}$ samples, the distribution modeling methods (i.e. OC-SVM, GMM, and NF) learn the underlying distribution of $\mathcal{F}_{\mathrm{monitor}}$ features and evaluate the likelihood of $\mathcal{F}_{\mathrm{test}}$ samples under this learned distribution. Although APS serves as a valuable baseline for comparative analysis, APS computational complexity makes it impractical for real-time deployment. In contrast, OC-SVM, GMM, and NF offer efficient real-time capable inference times suitable for operational deployment, as their distribution parameters are learned offline.
@@ D.1 Semantic Shift Experiments
APS OC-SVM GMM NF
AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95
CLIP RN50 0.85 0.82 0.49 0.85 0.83 0.51 0.90 0.88 0.36 0.82 0.78 0.52
RN101 0.83 0.82 0.55 0.83 0.82 0.65 0.87 0.85 0.46 0.79 0.74 0.60
RN50x4 0.84 0.81 0.55 0.84 0.81 0.62 0.88 0.86 0.42 0.84 0.79 0.50
RN50x16 0.88 0.86 0.44 0.89 0.88 0.39 0.89 0.87 0.38 0.83 0.80 0.51
RN50x64 0.94 0.93 0.24 0.94 0.93 0.24 0.94 0.93 0.23 0.88 0.84 0.37
ViT-B/32 0.86 0.84 0.51 0.86 0.85 0.53 0.88 0.85 0.38 0.87 0.84 0.50
ViT-B/16 0.88 0.86 0.38 0.89 0.88 0.43 0.91 0.89 0.33 0.91 0.88 0.34
ViT-L/14 0.87 0.85 0.45 0.89 0.87 0.41 0.85 0.83 0.47 0.88 0.86 0.41
DINO ViT-S/16 0.75 0.71 0.59 0.74 0.71 0.76 0.81 0.78 0.50 0.81 0.78 0.55
ViT-S/8 0.71 0.65 0.61 0.75 0.71 0.60 0.82 0.81 0.51 0.79 0.77 0.60
ViT-B/16 0.79 0.75 0.52 0.79 0.76 0.59 0.84 0.82 0.55 0.75 0.71 0.66
ViT-B/8 0.81 0.76 0.49 0.77 0.75 0.66 0.87 0.86 0.45 0.82 0.81 0.59
RN50 0.64 0.63 0.87 0.61 0.60 0.91 0.68 0.67 0.80 0.61 0.60 0.88
DINOv2 ViT-S/14 0.86 0.80 0.34 0.88 0.83 0.31 0.81 0.76 0.51 0.78 0.74 0.65
ViT-B/14 0.86 0.82 0.45 0.87 0.84 0.53 0.82 0.78 0.54 0.85 0.80 0.48
ViT-L/14 0.89 0.87 0.43 0.84 0.82 0.68 0.82 0.77 0.52 0.83 0.78 0.54
ViT-G/14 0.90 0.87 0.30 0.85 0.82 0.63 0.86 0.78 0.39 0.83 0.75 0.46
Grounding DINO SwinB 0.88 0.84 0.47 0.84 0.80 0.54 0.91 0.89 0.32 0.95 0.94 0.16
SwinT 0.85 0.80 0.53 0.83 0.79 0.59 0.92 0.90 0.28 0.96 0.95 0.16
@@
CLIP RN50 0.87 0.89 0.65 0.87 0.90 0.71 0.89 0.92 0.66 0.92 0.94 0.57
RN101 0.89 0.92 0.54 0.89 0.92 0.60 0.91 0.93 0.55 0.93 0.95 0.49
RN50x4 0.90 0.92 0.52 0.91 0.93 0.58 0.93 0.95 0.48 0.94 0.95 0.42
RN50x16 0.92 0.94 0.48 0.92 0.94 0.50 0.93 0.94 0.49 0.94 0.96 0.49
RN50x64 0.93 0.94 0.54 0.93 0.95 0.54 0.94 0.95 0.53 0.94 0.96 0.43
ViT-B/32 0.88 0.90 0.57 0.88 0.90 0.61 0.91 0.93 0.52 0.92 0.94 0.54
ViT-B/16 0.91 0.93 0.50 0.91 0.93 0.54 0.92 0.94 0.48 0.93 0.95 0.49
ViT-L/14 0.92 0.94 0.48 0.92 0.94 0.53 0.93 0.95 0.46 0.94 0.96 0.50
DINO ViT-S/16 0.79 0.80 0.74 0.81 0.83 0.74 0.85 0.87 0.66 0.88 0.91 0.63
ViT-S/8 0.83 0.85 0.69 0.84 0.87 0.76 0.87 0.90 0.61 0.89 0.91 0.64
ViT-B/16 0.81 0.82 0.69 0.83 0.85 0.75 0.86 0.88 0.60 0.84 0.87 0.80
ViT-B/8 0.82 0.84 0.64 0.83 0.86 0.70 0.86 0.89 0.59 0.88 0.91 0.65
RN50 0.55 0.54 0.92 0.50 0.50 0.92 0.55 0.54 0.94 0.58 0.57 0.89
DINOv2 ViT-S/14 0.75 0.77 0.77 0.75 0.78 0.82 0.79 0.82 0.77 0.85 0.88 0.74
ViT-B/14 0.79 0.80 0.71 0.80 0.82 0.73 0.84 0.86 0.64 0.87 0.89 0.67
ViT-L/14 0.78 0.79 0.77 0.78 0.81 0.80 0.82 0.85 0.74 0.86 0.88 0.63
ViT-G/14 0.79 0.80 0.73 0.81 0.83 0.69 0.82 0.85 0.71 0.86 0.89 0.69
Grounding DINO SwinB 0.60 0.56 0.83 0.58 0.54 0.85 0.76 0.68 0.58 0.99 0.99 0.02
SwinT 0.57 0.54 0.89 0.57 0.54 0.91 0.64 0.58 0.78 0.91 0.84 0.20
Table 9: VFM performance comparison across (i.e. ResNet, ViT, and Swin) architectures on Lost and Found [57] with different intensity values and Bravo-Synobj [49] datasets. AUROC↑, AUPR↑, and FPR95↓ metrics shown for APS, OC-SVM, GMM, and NF methods. Best values in bold.
@@
The extended analysis reveals interesting patterns in cross-dataset performance. Models that excel on real-world data don’t necessarily maintain their advantage on synthetic data. This is particularly evident in CLIP’s RN50x64 performance. which dominates on Lost and Found (AUROC: 0.94) but shows relatively lower performance on Bravo-Synobj (AUROC: 0.93). Conversely. some architectures like CLIP’s ViT-L/14 demonstrate more balanced performance across both datasets. suggesting better generalization capabilities.
@@
The full backbone comparison enables analysis of patch size effects in transformer-based architectures. In DINO models. smaller patch sizes (ViT-B/8 vs ViT-B/16) generally lead to improved performance. particularly evident in GMM results (AUROC: 0.87 vs 0.84 on Lost and Found). This suggests that finer-grained feature extraction benefits semantic shift detection. though at the cost of computational complexity.

### v3

@@ Results
APS MFS OC-SVM GMM NF
AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95
Trained on ImageNet RN50 0.73 0.69 0.70 0.73 0.69 0.70 0.76 0.71 0.62 0.70 0.67 0.78 0.75 0.71 0.63
RN101 0.70 0.64 0.68 0.70 0.64 0.68 0.72 0.65 0.60 0.69 0.65 0.81 0.75 0.70 0.63
RN152 0.68 0.63 0.74 0.68 0.63 0.74 0.72 0.68 0.77 0.67 0.62 0.77 0.70 0.65 0.73
ViT-B/16 0.79 0.75 0.53 0.79 0.75 0.53 0.83 0.81 0.53 0.81 0.78 0.53 0.82 0.79 0.56
ViT-L/16 0.88 0.85 0.39 0.88 0.85 0.39 0.86 0.83 0.47 0.88 0.85 0.40 0.88 0.85 0.42
SwinB 0.84 0.82 0.53 0.84 0.82 0.53 0.83 0.84 0.54 0.85 0.84 0.53 0.86 0.85 0.44
SwinT 0.72 0.64 0.62 0.72 0.64 0.62 0.74 0.70 0.75 0.73 0.66 0.65 0.72 0.67 0.77
Autoencoder 128 0.66 0.61 0.71 0.55 0.52 0.80 0.62 0.56 0.74 0.62 0.55 0.79 0.42 0.43 0.94
256 0.68 0.62 0.69 0.70 0.67 0.69 0.68 0.64 0.77 0.52 0.48 0.87 0.60 0.56 0.82
512 0.65 0.61 0.73 0.68 0.63 0.70 0.62 0.58 0.74 0.61 0.58 0.88 0.55 0.53 0.88
1024 0.70 0.67 0.74 0.77 0.74 0.56 0.62 0.57 0.84 0.65 0.60 0.83 0.68 0.63 0.65
Ours Clip-RN50 0.85 0.82 0.49 0.85 0.82 0.49 0.90 0.88 0.36 0.85 0.83 0.51 0.82 0.78 0.52
Clip-RN50x64 0.94 0.93 0.24 0.94 0.93 0.24 0.94 0.93 0.23 0.94 0.93 0.24 0.88 0.84 0.37
SwinB 0.88 0.84 0.47 0.88 0.84 0.47 0.91 0.89 0.32 0.84 0.80 0.54 0.95 0.94 0.16
@@
Trained on ImageNet RN50 0.90 0.89 0.33 0.79 0.83 0.82 0.94 0.93 0.23 0.92 0.91 0.27 0.97 0.97 0.12
RN101 0.89 0.87 0.36 0.81 0.85 0.79 0.93 0.92 0.27 0.91 0.90 0.30 0.97 0.96 0.12
RN152 0.90 0.89 0.35 0.80 0.84 0.78 0.95 0.94 0.18 0.93 0.91 0.29 0.97 0.96 0.10
ViT-B/16 0.86 0.82 0.38 0.81 0.81 0.69 0.93 0.92 0.24 0.91 0.88 0.26 0.97 0.96 0.12
ViT-L/16 0.88 0.84 0.31 0.85 0.88 0.70 0.92 0.91 0.24 0.91 0.88 0.23 0.98 0.97 0.07
SwinB 0.90 0.88 0.31 0.84 0.86 0.67 0.96 0.96 0.14 0.95 0.93 0.19 0.99 0.99 0.05
SwinT 0.88 0.84 0.31 0.76 0.75 0.77 0.95 0.93 0.16 0.93 0.90 0.20 0.99 0.98 0.04
Autoencoder 128 0.60 0.59 0.87 0.58 0.55 0.87 0.59 0.56 0.86 0.60 0.58 0.88 0.55 0.52 0.87
256 0.60 0.59 0.85 0.57 0.55 0.87 0.58 0.55 0.87 0.54 0.53 0.90 0.59 0.55 0.85
512 0.60 0.60 0.84 0.56 0.54 0.88 0.56 0.55 0.87 0.60 0.59 0.88 0.57 0.54 0.85
1024 0.57 0.55 0.87 0.59 0.56 0.86 0.60 0.57 0.86 0.52 0.51 0.86 0.63 0.59 0.81
Ours Clip-RN50 0.76 0.72 0.59 0.76 0.72 0.59 0.85 0.82 0.43 0.75 0.71 0.64 0.95 0.92 0.16
Clip-RN50x64 0.99 0.99 0.03 0.99 0.99 0.03 1.00 1.00 0.01 1.00 1.00 0.01 0.99 0.99 0.04
SwinB 1.00 1.00 0.02 1.00 1.00 0.02 1.00 1.00 0.00 1.00 0.99 0.02 1.00 1.00 0.00
@@
Table 6: Comparative evaluation of out-of-distribution detection performance using ImageNet-pretrained models, autoencoders, and VFMs on Lost and Found [59] and Bravo-Synrain [50] benchmarks. We evaluate using AUROC, AUPR, and FPR95 metrics across five methods: APS, MFS, OC-SVM, GMM, and NF. For VFMs, we utilize image encoders from CLIP [61] with ResNet-50 and ResNet-50x64 backbones, and from Grounding-DINO [45] with SwinB backbone. Best performances for each metric and dataset are highlighted in bold.
A.3 Real-time Capability of Chosen Models
@@
A critical consideration for practical deployment in AD systems is computational efficiency. Our approach leverages only the image encoder components of Vision Foundation Models (VFMs), specifically employing CLIP’s ResNet variants [61] and Grounding DINO’s Swin Transformer backbones [45]. Recent literature substantiates the real-time capabilities of these architectures:
@@
  • Zhao et al.[90] demonstrate that DETR-based architectures achieve real-time performance while surpassing YOLO-based models in detection accuracy.
@@
  • Kassab et al.[36] have successfully deployed CLIP’s ResNet backbones for real-time visual scene understanding in SLAM applications, validating their computational feasibility.
Our framework offers additional deployment flexibility through temporal and spatial optimization strategies, including:
  • Processing every $n$th frame to reduce computational load
  • Operating on downsampled inputs while maintaining monitoring efficacy
To quantify real-time performance, we benchmarked all feature extractors on hardware (Intel Xeon Gold 6326 CPU, NVIDIA A40 GPU) using randomly sampled 50 full-resolution (2048×1024) images from Cityscapes. Table 7 presents frames-per-second (FPS) rates with standard deviations across 5 measurement runs.
Model Backbone FPS
Mean Std
CLIP RN50 63.69 0.81
RN101 63.29 0.40
RN50x4 56.50 0.32
RN50x16 42.37 0.90
RN50x64 31.15 0.19
ViT-B/32 63.29 1.60
ViT-B/16 56.18 0.32
ViT-L/14 49.50 0.74
DINO ViT-S/16 59.17 2.10
ViT-S/8 49.50 0.00
ViT-B/16 49.02 3.37
ViT-B/8 42.74 0.18
RN50 56.50 0.00
DINOv2 ViT-S/14 53.19 1.70
ViT-B/14 57.14 2.94
ViT-L/14 47.62 0.23
ViT-G/14 35.34 0.87
Grounding DINO SwinB 16.31 0.08
SwinT 26.74 2.43
Table 7: Inference speed of VFMs’ feature extractors, measured in frames per second (FPS) per image on high-resolution 2048×1024 input images. The FPS values represent the mean and standard deviation (Std) across multiple runs.
Our results indicate that CLIP’s ResNet and ViT backbones provide the highest throughput, with RN50 and ViT-B/32 both exceeding 63 FPS. The DINOv2 and DINO variants maintain strong performance between 35-59 FPS. While Grounding DINO models demonstrate lower throughput (SwinB: 16.31 FPS, SwinT: 26.74 FPS), they remain viable for real-time applications—particularly given their superior OOD detection performance. These throughput rates can be improved by employing our proposed temporal optimization (processing every $n$th frame) or spatial optimization (downsampling input resolution). Recent work by Zhao et al. [90] has shown that DETR-based architectures can surpass YOLO [15] in the inference with their optimizations which can be used in our case for Grounding-DINO measurements.
For distribution modeling approaches, computational efficiency varies significantly. APS requires computing pairwise similarities between all monitor and test samples, making it computationally prohibitive for real-time applications. In contrast, MFS, OC-SVM, GMM, and NF all offer efficient inference after offline training, as they only require evaluating new samples against pre-computed distribution parameters. Among these approaches, MFS provides the fastest inference but with reduced detection performance, while GMM and NF offer optimal performance-efficiency trade-offs for operational deployment.
@@ D.1 Semantic Shift Experiments
APS MFS OC-SVM GMM NF
AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95
CLIP RN50 0.85 0.82 0.49 0.84 0.82 0.52 0.85 0.83 0.51 0.90 0.88 0.36 0.82 0.78 0.52
RN101 0.83 0.82 0.55 0.82 0.81 0.58 0.83 0.82 0.65 0.87 0.85 0.46 0.79 0.74 0.60
RN50x4 0.84 0.81 0.55 0.83 0.82 0.58 0.84 0.81 0.62 0.88 0.86 0.42 0.84 0.79 0.50
RN50x16 0.88 0.86 0.44 0.88 0.88 0.46 0.89 0.88 0.39 0.89 0.87 0.38 0.83 0.80 0.51
RN50x64 0.94 0.93 0.24 0.93 0.93 0.27 0.94 0.93 0.24 0.94 0.93 0.23 0.88 0.84 0.37
ViT-B/32 0.86 0.84 0.51 0.83 0.82 0.60 0.86 0.85 0.53 0.88 0.85 0.38 0.87 0.84 0.50
ViT-B/16 0.88 0.86 0.38 0.86 0.85 0.41 0.89 0.88 0.43 0.91 0.89 0.33 0.91 0.88 0.34
ViT-L/14 0.87 0.85 0.45 0.89 0.89 0.41 0.89 0.87 0.41 0.85 0.83 0.47 0.88 0.86 0.41
DINO ViT-S/16 0.75 0.71 0.59 0.71 0.63 0.60 0.74 0.71 0.76 0.81 0.78 0.50 0.81 0.78 0.55
ViT-S/8 0.71 0.65 0.61 0.66 0.59 0.66 0.75 0.71 0.60 0.82 0.81 0.51 0.79 0.77 0.60
ViT-B/16 0.79 0.75 0.52 0.72 0.64 0.60 0.79 0.76 0.59 0.84 0.82 0.55 0.75 0.71 0.66
ViT-B/8 0.81 0.76 0.49 0.73 0.64 0.59 0.77 0.75 0.66 0.87 0.86 0.45 0.82 0.81 0.59
RN50 0.64 0.63 0.87 0.65 0.62 0.83 0.61 0.60 0.91 0.68 0.67 0.80 0.61 0.60 0.88
DINOv2 ViT-S/14 0.86 0.80 0.34 0.86 0.80 0.36 0.88 0.83 0.31 0.81 0.76 0.51 0.78 0.74 0.65
ViT-B/14 0.86 0.82 0.45 0.86 0.82 0.44 0.87 0.84 0.53 0.82 0.78 0.54 0.85 0.80 0.48
ViT-L/14 0.89 0.87 0.43 0.91 0.88 0.37 0.84 0.82 0.68 0.82 0.77 0.52 0.83 0.78 0.54
ViT-G/14 0.90 0.87 0.30 0.89 0.86 0.35 0.85 0.82 0.63 0.86 0.78 0.39 0.83 0.75 0.46
Grounding DINO SwinB 0.88 0.84 0.47 0.82 0.76 0.52 0.84 0.80 0.54 0.91 0.89 0.32 0.95 0.94 0.16
SwinT 0.85 0.80 0.53 0.79 0.73 0.60 0.83 0.79 0.59 0.92 0.90 0.28 0.96 0.95 0.16
@@
CLIP RN50 0.87 0.89 0.65 0.87 0.89 0.65 0.87 0.90 0.71 0.89 0.92 0.66 0.92 0.94 0.57
RN101 0.89 0.92 0.54 0.89 0.92 0.54 0.89 0.92 0.60 0.91 0.93 0.55 0.93 0.95 0.49
RN50x4 0.90 0.92 0.52 0.90 0.92 0.52 0.91 0.93 0.58 0.93 0.95 0.48 0.94 0.95 0.42
RN50x16 0.92 0.94 0.48 0.92 0.94 0.49 0.92 0.94 0.50 0.93 0.94 0.49 0.94 0.96 0.49
RN50x64 0.93 0.94 0.54 0.93 0.94 0.54 0.93 0.95 0.54 0.94 0.95 0.53 0.94 0.96 0.43
ViT-B/32 0.88 0.90 0.57 0.88 0.90 0.56 0.88 0.90 0.61 0.91 0.93 0.52 0.92 0.94 0.54
ViT-B/16 0.91 0.93 0.50 0.91 0.93 0.50 0.91 0.93 0.54 0.92 0.94 0.48 0.93 0.95 0.49
ViT-L/14 0.92 0.94 0.48 0.92 0.94 0.48 0.92 0.94 0.53 0.93 0.95 0.46 0.94 0.96 0.50
DINO ViT-S/16 0.79 0.80 0.74 0.79 0.80 0.74 0.81 0.83 0.74 0.85 0.87 0.66 0.88 0.91 0.63
ViT-S/8 0.83 0.85 0.69 0.83 0.85 0.69 0.84 0.87 0.76 0.87 0.90 0.61 0.89 0.91 0.64
ViT-B/16 0.81 0.82 0.69 0.81 0.82 0.69 0.83 0.85 0.75 0.86 0.88 0.60 0.84 0.87 0.80
ViT-B/8 0.82 0.84 0.64 0.82 0.84 0.64 0.83 0.86 0.70 0.86 0.89 0.59 0.88 0.91 0.65
RN50 0.55 0.54 0.92 0.55 0.54 0.92 0.50 0.50 0.92 0.55 0.54 0.94 0.58 0.57 0.89
DINOv2 ViT-S/14 0.75 0.77 0.77 0.75 0.77 0.77 0.75 0.78 0.82 0.79 0.82 0.77 0.85 0.88 0.74
ViT-B/14 0.79 0.80 0.71 0.79 0.80 0.71 0.80 0.82 0.73 0.84 0.86 0.64 0.87 0.89 0.67
ViT-L/14 0.78 0.79 0.77 0.78 0.79 0.77 0.78 0.81 0.80 0.82 0.85 0.74 0.86 0.88 0.63
ViT-G/14 0.79 0.80 0.73 0.79 0.80 0.73 0.81 0.83 0.69 0.82 0.85 0.71 0.86 0.89 0.69
Grounding DINO SwinB 0.60 0.56 0.83 0.61 0.56 0.82 0.58 0.54 0.85 0.76 0.68 0.58 0.99 0.99 0.02
SwinT 0.57 0.54 0.89 0.58 0.54 0.90 0.57 0.54 0.91 0.64 0.58 0.78 0.91 0.84 0.20
Table 10: VFM performance comparison across (i.e. ResNet, ViT, and Swin) architectures on Lost and Found [59] with different intensity values and Bravo-Synobj [50] datasets. AUROC↑, AUPR↑, and FPR95↓ metrics shown for APS, MFS, OC-SVM, GMM, and NF methods. Best values in bold.
@@
The extended analysis reveals interesting patterns in cross-dataset performance. Models that excel on real-world data don’t necessarily maintain their advantage on synthetic data. This is particularly evident in CLIP’s RN50x64 performance, which dominates on Lost and Found (AUROC: 0.94) but shows relatively lower performance on Bravo-Synobj (AUROC: 0.93). Conversely. some architectures like CLIP’s ViT-L/14 demonstrate more balanced performance across both datasets. suggesting better generalization capabilities.
@@
The full backbone comparison enables analysis of patch size effects in transformer-based architectures. In DINO models, smaller patch sizes (ViT-B/8 vs ViT-B/16) generally lead to improved performance. Particularly evident in GMM results (AUROC: 0.87 vs 0.84 on Lost and Found). This suggests that finer-grained feature extraction benefits semantic shift detection. though at the cost of computational complexity.